# Sample Complexity of Composite Quantum Hypothesis Testing

### v2

@@ II-C Composite Quantum Hypothesis Testing
In the following sections, we provide the first study of the sample complexity of composite QHT. This analysis unveils how the sample complexity depends on key problem parameters, including the “similarity" of the uncertainty sets, the number of elements they contain, and the dimensionality of the underlying Hilbert space. For notational convenience, we abbreviate $n^{\ast}(p,\mathcal{D}_{1},\mathcal{D}_{2},\delta)$ (or $n^{\ast}(p,\rho_{1},\rho_{2},\delta)$) as $n^{\ast}(\delta)$ whenever the prior and the sets (or states) are clear from the context. Finally, we exclude the following trivial cases from our analysis, which generalizes the regimes identified in [8, Remark 2]. A proof can be found in Appendix -A.
@@ III-A Analysis of Min-max Expected Error Probability
where in the first equality, we used linearity of $P_{e}$ in its argument quantum states, whereby the supremum is achieved at the extreme points of the convex hull. The second equality follows from applying Sion’s minimax theorem [5, Lemma A.1] to swap the order of $\sup$ and $\inf$, since $P_{e}(\Pi,p,\sigma_{1,n},\sigma_{2,n})$ is linear in $\Pi$, and the $\mathcal{C}_{i,n}$ are closed, compact and convex sets. The result then follows from the last equality via (3). ∎
@@
The first inequality uses the well-known quantum Chernoff Bound [1, Theorem 1], while the second inequality follows by choosing $s=1/2$. The last inequality follows from [2, Eq. 28] and the definition of Uhlmann fidelity. ∎
@@ III-C2 Uncertainty Sets of Finite Cardinality
For $i=1,2$, let $\sigma_{i,n}\in\mathcal{C}_{i,n}$ denote quantum states in the respective convex hulls, which can be written equivalently as $\sigma_{1,n}=\sum_{j=1}^{m_{1}}p_{j}\rho_{1,j}^{\otimes n}$ and $\sigma_{2,n}=\sum_{k=1}^{m_{2}}q_{k}\rho_{2,k}^{\otimes n},$ where $\rho_{1,j}\in\mathcal{D}_{1}$, $\rho_{2,k}\in\mathcal{D}_{2}$ and $\{p_{j}\},\{q_{k}\}$ denote probability vectors. We then have
$F(\sigma_{1,n},\sigma_{2,n})$ $=\left\|\sqrt{\sum_{j=1}^{m_{1}}p_{j}\rho_{1,j}^{\otimes n}}\sqrt{\sum_{k=1}^{m_{2}}q_{k}\rho_{2,k}^{\otimes n}}\right\|_{1}$
$\leq\left\|\sum_{j=1}^{m_{1}}\sqrt{p_{j}\rho_{1,j}^{\otimes n}}\sum_{k=1}^{m_{2}}\sqrt{q_{k}\rho_{2,k}^{\otimes n}}\right\|_{1}$
$\leq\sup_{\rho_{i}\in\mathcal{D}_{i}}\left\|\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sqrt{\rho_{1}^{\otimes n}}\sqrt{\rho_{2}^{\otimes n}}\right\|_{1}$
$=\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sup_{\rho_{i}\in\mathcal{D}_{i}}\left\|\sqrt{\rho_{1}^{\otimes n}}\sqrt{\rho_{2}^{\otimes n}}\right\|_{1}$
@@
where the first inequality follows by using Rotfel’d trace inequality twice, since the square root is concave [17, Theorem 1.1] and the last inequality follows since $\sum_{j=1}^{m_{1}}\sqrt{p_{j}}\leq\sqrt{m_{1}}$. ∎
@@ A CPTP map $\mathcal{M}$ is $\varepsilon$-LDPQ for $\varepsilon\geq 0$ if
Using [7, Lemma 1], we have that $\ln F(\rho_{1},\rho_{2})\leq-\ln\left(1-\frac{1}{2}H_{1/2}(\rho_{1}\|\rho_{2})\right)\leq H_{1/2}(\rho_{1}\|\rho_{2})$, where we used $\ln(1-x)\leq 2x$ for all $x\in[0,\frac{1}{2}]$. Applying this to the lower bound of Theorem˜2, and using [7, (4.12)] gives n^∗_ε(δ,p,ρ_1,ρ_2) ≥(eε+1)lnp(1-p)δ2(eε-1)H1/2(ρ1ρ2). Additionally, [7, (4.25)] yields that n^∗_ε(δ,p,ρ_1,ρ_2) ≥(1-δ/p)2e-ε(eε-1)2E1(ρ1ρ2)2. The proof then follows from noting that $n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})$ for any $(\rho_{1},\rho_{2})\in\mathcal{D}_{1}\times\mathcal{D}_{2}$, whereby we get $n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq\sup_{\rho_{i}\in\mathcal{D}_{i}}n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})$. ∎
The above result directly generalizes the sample complexity of locally differentially private simple QHT, recovering [7, Theorems 4.6 and 4.9] for $p=0.5$, $\delta=0.1$ and $|\mathcal{D}_{i}|=1$ for $i=1,2$.
@@
Let $\mathcal{B}$ be the $\varepsilon$-LDP channel as constructed in the proof of [7, Theorem 4.2]. Then, we have
@@
where the first inequality uses $1-x\leq-\ln x$ for all $x\in[0,1]$, the second uses [7, Equation 4.7], and the last inequality uses the specific channel $\mathcal{B}$ to evaluate $E_{1}(\mathcal{B}(\rho_{1})\|\mathcal{B}(\rho_{2}))$. ∎

### v3

@@ II-C Composite Quantum Hypothesis Testing
In the following sections, we provide the first study of the sample complexity of composite QHT. This analysis unveils how the sample complexity depends on key problem parameters, including the “similarity" of the uncertainty sets, the number of elements they contain, and the dimensionality of the underlying Hilbert space. For notational convenience, we abbreviate $n^{\ast}(p,\mathcal{D}_{1},\mathcal{D}_{2},\delta)$ (or $n^{\ast}(p,\rho_{1},\rho_{2},\delta)$) as $n^{\ast}(\delta)$ whenever the prior and the sets (or states) are clear from the context. Finally, we exclude the following trivial cases from our analysis, which generalizes the regimes identified in [9, Remark 2]. A proof can be found in Appendix -A.
@@ III-A Analysis of Min-max Expected Error Probability
where in the first equality, we used linearity of $P_{e}$ in its argument quantum states, whereby the supremum is achieved at the extreme points of the convex hull. The second equality follows from applying Sion’s minimax theorem [6, Lemma A.1] to swap the order of $\sup$ and $\inf$, since $P_{e}(\Pi,p,\sigma_{1,n},\sigma_{2,n})$ is linear in $\Pi$, and the $\mathcal{C}_{i,n}$ are closed, compact and convex sets. The result then follows from the last equality via (3). ∎
@@
The first inequality uses the well-known quantum Chernoff Bound [1, Theorem 1], while the second inequality follows by choosing $s=1/2$. The last inequality follows from [3, Eq. 28] and the definition of Uhlmann fidelity. ∎
@@ III-C2 Uncertainty Sets of Finite Cardinality
For $i=1,2$, let $\sigma_{i,n}\in\mathcal{C}_{i,n}$ denote any quantum states in the respective convex hulls, which can be written equivalently as $\sigma_{1,n}=\sum_{j=1}^{m_{1}}p_{j}\rho_{1,j}^{\otimes n}$ and $\sigma_{2,n}=\sum_{k=1}^{m_{2}}q_{k}\rho_{2,k}^{\otimes n},$ where $\rho_{1,j}\in\mathcal{D}_{1}$, $\rho_{2,k}\in\mathcal{D}_{2}$ and $\{p_{j}\},\{q_{k}\}$ denote probability vectors. For any $\sigma_{i,n}\in\mathcal{C}_{i,n}$, the following set of relations hold,
$F(\sigma_{1,n},\sigma_{2,n})$ $=F\left(\sum_{j=1}^{m_{1}}p_{j}\rho_{1,j}^{\otimes n},\sum_{k=1}^{m_{2}}q_{k}\rho_{2,k}^{\otimes n}\right)$
$\leq\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}F(p_{j}\rho_{1,j}^{\otimes n},q_{k}\rho_{2,k}^{\otimes n})$
$=\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}F(\rho_{1,j},\rho_{2,k})^{n}$
$\leq\sum_{j=1}^{m_{1}}\sum_{k=1}^{m_{2}}\sqrt{p_{j}q_{k}}\sup_{\rho_{i}\in\mathcal{D}_{i}}F(\rho_{1},\rho_{2})^{n}$
@@
where the first inequality follows by sub-additivity of Fidelity [2, Lemma 4.9], and the last inequality follows by Cauchy-Schwarz: $\sum_{j=1}^{m_{1}}(\sqrt{p_{j}}\cdot 1)\leq\sqrt{(\sum_{j=1}^{m_{1}}p_{j})(\sum_{j=1}^{m_{1}}1^{2})}=\sqrt{m_{1}}$. ∎
@@ A CPTP map $\mathcal{M}$ is $\varepsilon$-LDPQ for $\varepsilon\geq 0$ if
Using [8, Lemma 1], we have that $\ln F(\rho_{1},\rho_{2})\leq-\ln\left(1-\frac{1}{2}H_{1/2}(\rho_{1}\|\rho_{2})\right)\leq H_{1/2}(\rho_{1}\|\rho_{2})$, where we used $\ln(1-x)\leq 2x$ for all $x\in[0,\frac{1}{2}]$. Applying this to the lower bound of Theorem˜2, and using [8, (4.12)] gives n^∗_ε(δ,p,ρ_1,ρ_2) ≥(eε+1)lnp(1-p)δ2(eε-1)H1/2(ρ1ρ2). Additionally, [8, (4.25)] yields that n^∗_ε(δ,p,ρ_1,ρ_2) ≥(1-δ/p)2e-ε(eε-1)2E1(ρ1ρ2)2. The proof then follows from noting that $n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})$ for any $(\rho_{1},\rho_{2})\in\mathcal{D}_{1}\times\mathcal{D}_{2}$, whereby we get $n^{\ast}_{\varepsilon}(\delta,p,\mathcal{D}_{1},\mathcal{D}_{2})\geq\sup_{\rho_{i}\in\mathcal{D}_{i}}n^{\ast}_{\varepsilon}(\delta,p,\rho_{1},\rho_{2})$. ∎
The above result directly generalizes the sample complexity of locally differentially private simple QHT, recovering [8, Theorems 4.6 and 4.9] for $p=0.5$, $\delta=0.1$ and $|\mathcal{D}_{i}|=1$ for $i=1,2$.
@@
Let $\mathcal{B}$ be the $\varepsilon$-LDP channel as constructed in the proof of [8, Theorem 4.2]. Then, we have
@@
where the first inequality uses $1-x\leq-\ln x$ for all $x\in[0,1]$, the second uses [8, Equation 4.7], and the last inequality uses the specific channel $\mathcal{B}$ to evaluate $E_{1}(\mathcal{B}(\rho_{1})\|\mathcal{B}(\rho_{2}))$. ∎